OpenStudy (anonymous):

Two cars are driving toward each other from opposite directions. If one is driving 55 mph, the other is driving 60 mph and the cars are 500 miles apart, how long will it be before they pass each other? Round to the nearest tenth of an hour. A. 4.3 hours B. 4.5 hours C. 5.1 hours D. 4.8 hours

OpenStudy (inkyvoyd):

@Joshuam1 , you do understand it is against the conditions of this site to use this site to cheat on your test?

OpenStudy (anonymous):

Yes

OpenStudy (inkyvoyd):

Well, I would suggest that you stop.

OpenStudy (anonymous):

I'm not cheating on a "test"

OpenStudy (inkyvoyd):

Doesn't matter. If it's homework, you are still cheating.

OpenStudy (anonymous):

I know what cheating is and I'll be sure not to do it

OpenStudy (inkyvoyd):

You previous closed questions suggest the contrary.

OpenStudy (anonymous):

Yeah it does seem that way.

OpenStudy (anonymous):

Thank you for reminding me Inkyvoyd

OpenStudy (anonymous):

T(55 + 60) = 500 T(115) = 500 T = 500/115 T = 4.34 hours Rounding to the nearest tenth gives 4.3 So it was A.